Advanced Modular Technology​ (AMT) makes energy​ cleaner, safer, more secure and more efficient. It typically exhibits net annual revenues that increase over a fairly long period. In the long​ run, an AMT project may be profitable as measured by​ IRR, but its simple payback period may be unacceptable. Evaluate this AMT project using the IRR method when the company MARR IS 12​% per year and its maximum allowable payback period is three years. What is your​ recommendation?Capital investment at time 0 Net revenues in year k $103,000 $22,000 + $9,000. (k-1) $9,000 5 years Market (salvage) value Li
Answer #1

IRR Criteria

Let IRR be I%, then

22000 * (P/A, I%,5) + 9000 * (P/G, I%,5) + 9000 * (P/F, I%,5) = 103000

Dividing by 1000

22 * (P/A, I%,5) + 9 * (P/G, I%,5) + 9 * (P/F, I%,5) = 103

using trail and error method

When I = 20%, Value of 22 * (P/A, I%,5) + 9 * (P/G, I%,5) + 9 * (P/F, I%,5) = 113.5654

When I = 23% , Value of 22 * (P/A, I%,5) + 9 * (P/G, I%,5) + 9 * (P/F, I%,5) = 105.0784

When I = 24% , Value of 22 * (P/A, I%,5) + 9 * (P/G, I%,5) + 9 * (P/F, I%,5) = 102.4626

using interpolation

i= 23% + (105.0784 - 103)/(105.0784 - 102.4626)*(24% - 23%)

i = 23% + 0.79%

i = 23.79%

As IRR> MARR, project is acceptable

Payback criteria

Cumulative cash flow in yr 0 = -103000

Cumulative cash flow in yr 1 = -103000 + 22000 = -81000

Cumulative cash flow in yr 2 = -81000 + 22000 + 9000 = -50000

Cumulative cash flow in yr 3 = -50000 + 22000 + 9000 + 9000 = -10000

Cumulative cash flow in yr 4 = -10000 + 22000 + 9000 + 9000 + 9000 = 39000

Simple payback period = 3 + 10000 / 49000 = 3.2 yrs

As payback period is more than 3 yrs, project is not acceptable